# Determining the optimal structural resolution of proteins through an information-theoretic analysis of their conformational ensemble

### v5

@@ 3.2 Simulation and analysis protocols
All simulations were performed with the Gromacs 2018 software 39. The protein topology was defined through the AMBER99SB-ILDN force field 40, and the TIP3P model was employed for the description of water molecules 41. Sodium and chloride ions were added at the physiological concentration of $150$ mM, neutralising the global electric charge in the simulation box. After energy minimisation, NVT and NPT equilibrations were performed using the velocity-rescale thermostat 42 and the Parrinello–Rahman barostat 43. A cut-off of $1.2$ nm was used for van der Waals interaction and for the short-range component of the Coulomb one. The long-range component of the Coulomb force, instead, was computed with the Particle Mesh Ewald algorithm. The LINCS 44 algorithm was employed to define the constraints on the hydrogen-containing bonds. An integration time step of $2$ fs was used. For each protein of the dataset, a trajectory of $500$ ns was collected; of these, the first $100$ ns are considered as equilibration and therefore discarded.
Analysis of the trajectories was performed with Gromacs tools and in-house scripts. Protein residue fluctuations were calculated with gmx rmsf, while the clustering and the identification of the representative conformations was performed with gmx cluster. The rendering of protein structures was performed with VMD 45 and with TCL scripts.
@@ 3.5 Mapping entropy
Mapping entropy quantifies the intrinsic loss of information that arises when a fine-grained system is projected onto a CG representation. This concept was first introduced within the framework of the relative entropy method 46 and has since been widely applied to the analysis of proteins as well as other systems 21, 28, 20, 47.
@@
The mapping entropy $S_{\mathrm{map}}$ is defined as the Kullback–Leibler divergence between the original microscopic distribution $p(\mathbf{x})$ and the reconstructed distribution $\tilde{p}(\mathbf{x})$ inferred from the coarse-grained representation:
@@
Mapping entropy calculations were performed using the EXCOGITO software package 48, a general-purpose toolkit for coarse-graining analysis. In our case, the mapping entropy was estimated as the Kullback–Leibler divergence between the full microscopic distribution and the reconstructed one, with structural similarity assessed through clustering based on the root square deviation (RSD), adopting the same distance threshold as in the PROPRE protocol. Further details on the methodology can be found in 48, 28, 46, 49.
@@ 4.1 Optimal number of sites vs number of protein residues
The PROPRE algorithm allows the identification of the optimal resolution for a specific protein, where the optimal resolution is intended as the fraction of the total number of sites (namely, of the protein heavy atoms) that has to be retained in order to minimize the information loss upon coarse-graining. It is natural to think that the optimal number of sites $N_{\mathrm{OPT}}$ depends on the specific system under investigation and its features, e.g. the protein size, function, and flexibility. To assess the possible presence of a general trend, we fist applied the PROPRE pipeline to $500$ ns-long MD trajectories of a set of $11$ different proteins (see Fig. 4), ranging from a molecular weight of $7$ kDa (1IGD) to $66$ kDa (1XWL). The systems were selected from a larger protein dataset in order to maximize the differences in tertiary structure and dynamical features (see Methods section). Despite the large differences in structure, size and dynamics, the analysis on the trajectories of the 11 proteins of the dataset suggests a direct link between the optimal number of retained sites ($N_{\mathrm{OPT}}$) and the total number of protein residues ($N_{\mathrm{AA}}$). Fig 5 (a,b) reports this relationship for two distinct strategies to sample the trajectory: striding, that is, retaining $1000$ equally spaced frames; and clustering, i.e. employing the $1000$ centroids of an UPGMA clustering. In both cases, a clear linear trend emerges, with correlation coefficients $R=0.99$ for both uniform and centroid sampling. The fitted slope of about $3.8$ in both cases reveals a systematic trend, whereby optimal coarse-graining retains almost four heavy atoms per residue, suggesting that a description of backbone plus minimal side chain captures the essential degrees of freedom.
Figure 5: (a,b) Scatter plots illustrating the number of optimal sites identified by PROPRE for each protein in relation to the number of residues. (a) Analysis based on 1000 equidistant frames (striding). (b) Analysis based on the centroids of 1000 clusters obtained from a UPGMA clustering procedure. Error bars indicate the standard deviation in the number of optimal sites within the most probable bin, as determined by PROPRE using the density-based protocol. (c) Box plots of the ratio between the number of optimal sites and the average sampled volume ($N_{\mathrm{OPT}}/V$) for the two strategies: striding (STR) and clustering (CLU). For STR, orange points indicate values for the AKE protein computed from the full trajectory, as well as separately for the open and closed conformations.
@@
Figure 6: Scatter plot of the ratio between optimal sites and residue number ($N_{\mathrm{OPT}}/N_{\mathrm{AA}}$) identified for adenylate kinase from the full trajectory and from the subsets corresponding to the open and closed states, as a function of the average radius of gyration sampled in each case. The dashed horizontal line indicates the reference value fitted on the 11-protein dataset.
@@
These results indicate that $N_{\mathrm{OPT}}$ decreases as the average radius of gyration increases (see Fig. 6). In other words, conformational ensembles spanning larger configurational regions (e.g., the full trajectory or the open branch) require fewer atoms to be retained, whereas subsets confined to smaller fluctuations (e.g., the closed branch) need more atoms to achieve the same information balance. This inverse relation is consistent with the notion, widely discussed in the context of dimensionality reduction and collective variable selection, that large-scale, highly collective motions can be described with a reduced number of degrees of freedom, while small-amplitude local fluctuations require higher resolution to be meaningfully distinguished 50, 51, 52.
In the specific case of AKE, previous work has shown that its hallmark open-to-closed transition is dominated by hinge-bending of the LID and NMP domains, a motion that is intrinsically low-dimensional 53. Our findings quantitatively support this perspective: the broader the conformational exploration, the fewer atomic degrees of freedom are needed for an informative representation. In contrast, restricted conformational subsets entail finer local rearrangements, thus requiring higher resolution to preserve structural variability.
@@
Figure 7: (a) Correlation between the elbow position ($N_{\mathrm{CG}}^{E}$) in the variance of the coarse-grained covariance ($\sigma^{2}_{\mathrm{cov}}$) and the optimal number of retained atoms $N_{\mathrm{OPT}}$ identified by PROPRE, across a dataset of 11 proteins. Blue markers correspond to the information–theoretic estimate $N_{\mathrm{OPT}}^{\mathrm{IT}}$ (slope $\mu=-1$), red markers to the relevance–maximum estimate $N_{\mathrm{OPT}}^{\mathrm{MR}}$ (slope $\mu=0$). The legend reports the Pearson correlation coefficient for each criterion, and error bars indicate the standard deviation of $N_{\mathrm{OPT}}$ within the most-probable bin returned by PROPRE. (b–c) Analysis of protein 1IGD. In both panels, vertical blue dashed line marks the elbow position ($N_{\mathrm{CG}}^{E}$), vertical dashed red lines delimit the core PROPRE interval, and the shaded region extends it by one standard deviation. (b) Distribution of the normalized coarse-grained covariance $\mathrm{cov}$ over 50 random atom selections for each $N_{\mathrm{CG}}$ ; color encodes the pointwise density, the solid red line traces the running average. (c) Empirical variance $\sigma^{2}_{\mathrm{cov}}(N_{\mathrm{CG}})$ (gray dots) and fitted power-law decay (red curve).
@@
Figure 8: (a) Correlation between the peak position in the variance of the mapping entropy ($N_{\mathrm{CG}}^{P}$) and the optimal number of retained atoms $N_{\mathrm{OPT}}$ identified by PROPRE, across a dataset of 11 proteins. Blue markers correspond to the information–theoretic estimate $N_{\mathrm{OPT}}^{\mathrm{IT}}$ (slope $\mu=-1$), red markers to the relevance–maximum estimate $N_{\mathrm{OPT}}^{\mathrm{MR}}$ (slope $\mu=0$). The legend reports the Pearson correlation coefficient for each criterion, and error bars indicate the standard deviation of $N_{\mathrm{OPT}}$ within the most-probable bin returned by PROPRE. (b–c) Analysis of protein 1IGD. In both panels, the vertical blue dashed line marks the peak position $N_{\mathrm{CG}}^{P}$, vertical dashed red lines delimit the core PROPRE interval, and the shaded region extends it by one standard deviation. (b) Distribution of the mapping entropy $S_{\mathrm{map}}$ over 50 random atom selections for each $N_{\mathrm{CG}}$; color encodes the pointwise density. The yellow and green triangles highlight the values of $S_{\mathrm{map}}$ obtained from the specific selections of all Cα atoms and all backbone atoms, respectively. (c) Empirical variance $\sigma^{2}_{\mathrm{Smap}}(N_{\mathrm{CG}})$ (gray dots) and its Gaussian-smoothed estimate $\mathcal{S}^{2}_{\mathrm{Smap}}$ (black line).
@@
The average value of $S_{\mathrm{map}}$ decreases smoothly as $N_{\mathrm{CG}}$ increases, without exhibiting a distinct elbow or plateau. This trend indicates that, on average, adding more atoms progressively improves the reconstruction quality, but no sharp resolution threshold emerges from the mean behaviour. It is worth noting that when $S_{\mathrm{map}}$ is computed for specific selections—namely all Cα atoms or all backbone atoms—the resulting values lie well above those obtained from random heavy atom subsets of the same size (see Fig. 8.b, yellow triangle for Cα and green triangle for backbone). This outcome indicates that such selections are not particularly informative for this dataset, as the proteins considered do not undergo large conformational rearrangements but instead sample local basins around their native globular structures, a type of exploration that is not well captured by Cα or backbone coordinates alone.
@@
By contrast, the variance of $S_{\mathrm{map}}$ across random selections—denoted $\sigma^{2}_{\mathrm{Smap}}$—exhibits a well-defined maximum. As shown in the lower-right panel of Fig. 8, $\sigma^{2}_{\mathrm{Smap}}$ is low at both extremes of $N_{\mathrm{CG}}$ for opposite reasons: at very low $N_{\mathrm{CG}}$, the number of retained atoms is too small to capture relevant information, so all mappings yield similarly high $S_{\mathrm{map}}$ values; at high $N_{\mathrm{CG}}$, nearly all atoms are retained, and the mappings perform equivalently well, leading to low $S_{\mathrm{map}}$. Between these limits, a distinct peak emerges, pointing to an intermediate resolution level at which information retention becomes highly sensitive to the specific atoms selected for the coarse-grained representation, so that alternative subsets of the same size can vary substantially in how well they capture the original system properties. In this regime, the large variance of $S_{\mathrm{map}}$ indicates that the amount of information preserved depends strongly on which atoms are chosen. When $N_{\mathrm{CG}}$ is fixed well below $n_{\mathrm{ha}}$, targeted optimisation that minimises $S_{\mathrm{map}}$ by selecting which atoms to retain enojys the greatest opportunity to reduce information loss and thereby enhance the coarse-grained representation without increasing resolution.
@@
To robustly detect the location of this maximum, the raw variance profile $\sigma^{2}_{\mathrm{Smap}}$ was smoothed using a Gaussian filter 54 with a standard deviation of 5. The resulting curve, denoted $\mathcal{S}^{2}_{\mathrm{Smap}}$, enables reliable identification of the peak resolution, labelled $N_{\mathrm{CG}}^{P}$.
@@
The panels on the right of Fig. 8 display the results for one representative protein (1IGD), showing both the variance profile and the resolution range identified by PROPRE. The vertical dashed red lines indicate the core PROPRE interval, defined between the number of atoms maximizing relevance and the point where the slope of the relevance–resolution curve equals 1-1. The surrounding shaded band represents an extended interval that incorporates uncertainty—specifically, one standard deviation below the relevance maximum and above the slope 1-1 point. Notably, the peak in $\mathcal{S}^{2}_{\mathrm{Smap}}$ lies close to this shaded region, suggesting good consistency between the statistical properties of $S_{\mathrm{map}}$ and the information-theoretic resolution inferred by PROPRE. SI, Sec.V provides an extended presentation of the results for each individual protein in the dataset.
@@ 5 Conclusions
More specifically, for the 11-protein dataset, a clear linear correlation emerges between system size and optimal resolution: the number of atoms to be retained increases proportionally with the number of residues. The slope of this correlation, close to four heavy atoms per amino acid, indicates the average amount of detail needed to preserve the essential information in the ensemble. The linearity of this relation is far from trivial: although it can be expected that larger proteins naturally require more atoms, there was no guarantee that there would have been a clear dependence, let alone a linear one. Moreover, this trend is particularly meaningful because it mirrors the average level of resolution adopted in widely used coarse-grained models, such as MARTINI 55 and SIRAH 56, thus offering a quantitative, data-driven basis for their empirical success. In MARTINI, four heavy atoms are typically mapped to a single coarse-grained bead, which for proteins translates to one backbone bead plus up to 4 side-chain beads per residue, depending on the size and polarity of the side chain. In SIRAH, a slightly finer resolution is employed, where 2 to 4 heavy atoms are typically mapped onto one interaction site. These intermediate levels of reduction have been shown to preserve the essential conformational space accessible to proteins, offering a good trade-off between computational speed and structural/thermodynamical accuracy over extended time and length scales. Importantly, a further reduction in resolution, below the four-to-one mapping, has been reported to compromise structural fidelity and the ability to capture key dynamical characteristics 57, 58. For the proteins in the dataset, the optimal number of sites predicted by the PROPRE pipeline is intermediate between the number of CG beads obtained from the MARTINI and SIRAH mappings (see SI, sec III). Our finding that the data-driven fit identifies an effective scale similar to that of these CG models thus supports the view that $\sim 4$ atoms per residue represents a natural balance point: it reduces complexity substantially while retaining the conformational ensemble characteristic of the atomistic system.
@@
RP acknowledges support from the Italian Ministry of Education, University and Research (MIUR) through the FARE grant for the project HAMMOCK (grant R18ZHWY3NC). RP acknowledges support from ICSC - Centro Nazionale di Ricerca in HPC, Big Data and Quantum Computing, funded by the European Union under NextGenerationEU. Views and opinions expressed are however those of the author(s) only and do not necessarily reflect those of the European Union or The European Research Executive Agency. Neither the European Union nor the granting authority can be held responsible for them.
@@
Additional methodological details, full parameter lists, and supplementary analyses including relevance–resolution curve processing, PROPRE outputs, covariance and mapping entropy analyses, and comparison with coarse-grained models (PDF).

### v6

@@ 3.2 Simulation and analysis protocols
All simulations were performed with the Gromacs 2018 software 41. The protein topology was defined through the AMBER99SB-ILDN force field 42, and the TIP3P model was employed for the description of water molecules 43. Sodium and chloride ions were added at the physiological concentration of $150$ mM, neutralising the global electric charge in the simulation box. After energy minimisation, NVT and NPT equilibrations were performed using the velocity-rescale thermostat 44 and the Parrinello-Rahman barostat 45. A cut-off of $1.2$ nm was used for van der Waals interaction and for the short-range component of the Coulomb one. The long-range component of the Coulomb force, instead, was computed with the Particle Mesh Ewald algorithm. The LINCS 46 algorithm was employed to define the constraints on the hydrogen-containing bonds. An integration time step of $2$ fs was used. For each protein of the dataset, a trajectory of $500$ ns was collected; of these, the first $100$ ns are considered as equilibration and therefore discarded.
Analysis of the trajectories was performed with Gromacs tools and in-house scripts. Protein residue fluctuations were calculated with gmx rmsf, while the clustering and the identification of the representative conformations was performed with gmx cluster. The rendering of protein structures was performed with VMD 47 and with TCL scripts.
@@ 3.5 Mapping entropy
Mapping entropy quantifies the intrinsic loss of information that arises when a fine-grained system is projected onto a CG representation. This concept was first introduced within the framework of the relative entropy method 48 and has since been widely applied to the analysis of proteins as well as other systems 23, 30, 22, 49.
@@
The mapping entropy $S_{\mathrm{map}}$ is defined as the Kullback-Leibler divergence between the original microscopic distribution $p(\mathbf{x})$ and the reconstructed distribution $\tilde{p}(\mathbf{x})$ inferred from the coarse-grained representation:
@@
Mapping entropy calculations were performed using the EXCOGITO software package 50, a general-purpose toolkit for coarse-graining analysis. In our case, the mapping entropy was estimated as the Kullback-Leibler divergence between the full microscopic distribution and the reconstructed one, with structural similarity assessed through clustering based on the root square deviation (RSD), adopting the same distance threshold as in the PROPRE protocol. Further details on the methodology can be found in 50, 30, 48, 51.
@@ 4.1 Optimal number of sites vs number of protein residues
The PROPRE algorithm allows the identification of the optimal resolution for a specific protein, where the optimal resolution is intended as the fraction of the total number of sites (namely, of the protein heavy atoms) that has to be retained in order to minimize the information loss upon coarse-graining. In the following, the optimal number of retained sites is the one that corresponds to slope 1-1 in the resolution-relevance curve; analogous results obtained maximising the relevance are reported in the SI.
It is natural to think that the optimal number of sites $N_{\mathrm{OPT}}$ depends on the specific system under investigation and its features, e.g. the protein size, function, and flexibility. To assess the possible presence of a general trend, we fist applied the PROPRE pipeline to $500$ ns-long MD trajectories of a set of $11$ different proteins (see Fig. 4), ranging from a molecular weight of $7$ kDa (1IGD) to $66$ kDa (1XWL). The systems were selected from a larger protein dataset in order to maximize the differences in tertiary structure and dynamical features (see Methods section). Despite the large differences in structure, size and dynamics, the analysis on the trajectories of the 11 proteins of the dataset suggests a direct link between the optimal number of retained sites ($N_{\mathrm{OPT}}$) and the total number of protein residues ($N_{\mathrm{AA}}$). Fig 5 (a,b) reports this relationship for two distinct strategies to sample the trajectory: striding, that is, retaining $1000$ equally spaced frames; and clustering, i.e. employing the $1000$ centroids of an UPGMA clustering. In both cases, a clear linear trend emerges, with Pearson correlation coefficients $R=0.99$ for both uniform and centroid sampling. The fitted slope of about $3.8$ in both cases reveals a systematic trend, whereby optimal coarse-graining retains almost four heavy atoms per residue, suggesting that a description of backbone plus minimal side chain captures the essential degrees of freedom.
Figure 5: (a,b) Scatter plots illustrating the number of optimal sites identified by PROPRE for each protein in relation to the number of residues; for both cases (striding and clustering) the correlation coefficient of the linear fit ($R^{2}$) is $0.98$, the Pearson correlation coefficient is $0.99$. (a) Analysis based on $1000$ equidistant frames extracted through striding. (b) Analysis based on the centroids of $1000$ clusters obtained from a UPGMA clustering procedure. Error bars indicate the standard deviation in the number of optimal sites within the most probable bin, as determined by PROPRE using the density-based protocol. (c) Box plots of the ratio between the number of optimal sites and the average sampled volume ($N_{\mathrm{OPT}}/V$) for the two strategies: striding (STR) and clustering (CLU). For STR, orange points indicate values for the AKE protein computed from the full trajectory, as well as separately for the open and closed conformations.
@@
Figure 6: Scatter plot of the ratio between the optimal site number and residue number ($N_{\mathrm{OPT}}/N_{\mathrm{AA}}$) identified for adenylate kinase from the full trajectory and from the subsets corresponding to the open and closed states, as a function of the average radius of gyration sampled in each case. The optimal atom number corresponds to the resolution point of slope 1-1. The dashed horizontal line indicates the reference value fitted on the 11-protein dataset.
@@
These results indicate that $N_{\mathrm{OPT}}$ decreases as the average radius of gyration increases (see Fig. 6). In other words, conformational ensembles spanning larger configurational regions (e.g., the full trajectory or the open branch) require fewer atoms to be retained, whereas subsets confined to smaller fluctuations (e.g., the closed branch) need more atoms to achieve the same information balance. This inverse relation is consistent with the notion, widely discussed in the context of dimensionality reduction and collective variable selection, that large-scale, highly collective motions can be described with a reduced number of degrees of freedom, while small-amplitude local fluctuations require higher resolution to be meaningfully distinguished 52, 53, 54.
In the specific case of AKE, previous work has shown that its hallmark open-to-closed transition is dominated by hinge-bending of the LID and NMP domains, a motion that is intrinsically low-dimensional 55. Our findings quantitatively support this perspective: the broader the conformational exploration, the fewer atomic degrees of freedom are needed for an informative representation. In contrast, restricted conformational subsets entail finer local rearrangements, thus requiring higher resolution to preserve structural variability.
@@
Figure 7: (a) Correlation between the elbow position ($N_{\mathrm{CG}}^{E}$) in the variance of the coarse-grained covariance ($\sigma^{2}_{\mathrm{cov}}$) and the optimal number of retained atoms $N_{\mathrm{OPT}}$ identified by PROPRE, across a dataset of 11 proteins. Blue markers correspond to the information-theoretic estimate $N_{\mathrm{OPT}}^{\mathrm{IT}}$ (slope $\mu=-1$), red markers to the relevance-maximum estimate $N_{\mathrm{OPT}}^{\mathrm{MR}}$ (slope $\mu=0$). The legend reports the Pearson correlation coefficient for each criterion, and error bars indicate the standard deviation of $N_{\mathrm{OPT}}$ within the most-probable bin returned by PROPRE. (b-c) Analysis of protein 1IGD. In both panels, vertical blue dashed line marks the elbow position ($N_{\mathrm{CG}}^{E}$), vertical dashed red lines delimit the core PROPRE interval, and the shaded region extends it by one standard deviation. (b) Distribution of the normalized coarse-grained covariance $\mathrm{cov}$ over 50 random atom selections for each $N_{\mathrm{CG}}$ ; color encodes the pointwise density, the solid red line traces the running average. (c) Empirical variance $\sigma^{2}_{\mathrm{cov}}(N_{\mathrm{CG}})$ (gray dots) and fitted power-law decay (red curve).
@@
Figure 8: (a) Correlation between the peak position in the variance of the mapping entropy ($N_{\mathrm{CG}}^{P}$) and the optimal number of retained atoms $N_{\mathrm{OPT}}$ identified by PROPRE, across a dataset of 11 proteins. Blue markers correspond to the information-theoretic estimate $N_{\mathrm{OPT}}^{\mathrm{IT}}$ (slope $\mu=-1$), red markers to the relevance-maximum estimate $N_{\mathrm{OPT}}^{\mathrm{MR}}$ (slope $\mu=0$). The legend reports the Pearson correlation coefficient for each criterion, and error bars indicate the standard deviation of $N_{\mathrm{OPT}}$ within the most-probable bin returned by PROPRE. (b-c) Analysis of protein 1IGD. In both panels, the vertical blue dashed line marks the peak position $N_{\mathrm{CG}}^{P}$, vertical dashed red lines delimit the core PROPRE interval, and the shaded region extends it by one standard deviation. (b) Distribution of the mapping entropy $S_{\mathrm{map}}$ over 50 random atom selections for each $N_{\mathrm{CG}}$; color encodes the pointwise density. The yellow and green triangles highlight the values of $S_{\mathrm{map}}$ obtained from the specific selections of all Cα atoms and all backbone atoms, respectively. (c) Empirical variance $\sigma^{2}_{\mathrm{Smap}}(N_{\mathrm{CG}})$ (gray dots) and its Gaussian-smoothed estimate $\mathcal{S}^{2}_{\mathrm{Smap}}$ (black line).
@@
The average value of $S_{\mathrm{map}}$ decreases smoothly as $N_{\mathrm{CG}}$ increases, without exhibiting a distinct elbow or plateau. This trend indicates that, on average, adding more atoms progressively improves the reconstruction quality, but no sharp resolution threshold emerges from the mean behaviour. It is worth noting that when $S_{\mathrm{map}}$ is computed for specific selections–namely all Cα atoms or all backbone atoms—the resulting values lie well above those obtained from random heavy atom subsets of the same size (see Fig. 8.b, yellow triangle for Cα and green triangle for backbone). This outcome indicates that such selections are not particularly informative for this dataset, as the proteins considered do not undergo large conformational rearrangements but instead sample local basins around their native globular structures, a type of exploration that is not well captured by Cα or backbone coordinates alone.
@@
By contrast, the variance of $S_{\mathrm{map}}$ across random selections–denoted $\sigma^{2}_{\mathrm{Smap}}$–exhibits a well-defined maximum. As shown in the lower-right panel of Fig. 8, $\sigma^{2}_{\mathrm{Smap}}$ is low at both extremes of $N_{\mathrm{CG}}$ for opposite reasons: at very low $N_{\mathrm{CG}}$, the number of retained atoms is too small to capture relevant information, so all mappings yield similarly high $S_{\mathrm{map}}$ values; at high $N_{\mathrm{CG}}$, nearly all atoms are retained, and the mappings perform equivalently well, leading to low $S_{\mathrm{map}}$. Between these limits, a distinct peak emerges, pointing to an intermediate resolution level at which information retention becomes highly sensitive to the specific atoms selected for the coarse-grained representation, so that alternative subsets of the same size can vary substantially in how well they capture the original system properties. In this regime, the large variance of $S_{\mathrm{map}}$ indicates that the amount of information preserved depends strongly on which atoms are chosen. When $N_{\mathrm{CG}}$ is fixed well below $n_{\mathrm{ha}}$, targeted optimisation that minimises $S_{\mathrm{map}}$ by selecting which atoms to retain enjoys the greatest opportunity to reduce information loss and thereby enhance the coarse-grained representation without increasing resolution.
@@
To robustly detect the location of this maximum, the raw variance profile $\sigma^{2}_{\mathrm{Smap}}$ was smoothed using a Gaussian filter 56 with a standard deviation of 5. The resulting curve, denoted $\mathcal{S}^{2}_{\mathrm{Smap}}$, enables reliable identification of the peak resolution, labelled $N_{\mathrm{CG}}^{P}$.
@@
The panels on the right of Fig. 8 display the results for one representative protein (1IGD), showing both the variance profile and the resolution range identified by PROPRE. The vertical dashed red lines indicate the core PROPRE interval, defined between the number of atoms maximizing relevance and the point where the slope of the relevance-resolution curve equals 1-1. The surrounding shaded band represents an extended interval that incorporates uncertainty–specifically, one standard deviation below the relevance maximum and above the slope 1-1 point. Notably, the peak in $\mathcal{S}^{2}_{\mathrm{Smap}}$ lies close to this shaded region, suggesting good consistency between the statistical properties of $S_{\mathrm{map}}$ and the information-theoretic resolution inferred by PROPRE. SI, Sec.V provides an extended presentation of the results for each individual protein in the dataset.
@@ 5 Conclusions
More specifically, for the 11-protein dataset, a clear linear correlation emerges between system size and optimal resolution: the number of atoms to be retained increases proportionally with the number of residues. The slope of this correlation, close to four heavy atoms per amino acid, indicates the average amount of detail needed to preserve the essential information in the ensemble. The linearity of this relation is far from trivial: although it can be expected that larger proteins naturally require more atoms, there was no guarantee that there would have been a clear dependence, let alone a linear one. Moreover, this trend is particularly meaningful because it mirrors the average level of resolution adopted in widely used coarse-grained models, such as MARTINI 57 and SIRAH 58, thus offering a quantitative, data-driven basis for their empirical success. In MARTINI, four heavy atoms are typically mapped to a single coarse-grained bead, which for proteins translates to one backbone bead plus up to 4 side-chain beads per residue, depending on the size and polarity of the side chain. In SIRAH, a slightly finer resolution is employed, where 2 to 4 heavy atoms are typically mapped onto one interaction site. These intermediate levels of reduction have been shown to preserve the essential conformational space accessible to proteins, offering a good trade-off between computational speed and structural/thermodynamical accuracy over extended time and length scales. Importantly, a further reduction in resolution, below the four-to-one mapping, has been reported to compromise structural fidelity and the ability to capture key dynamical characteristics 59, 60. For the proteins in the dataset, the optimal number of sites predicted by the PROPRE pipeline is intermediate between the number of CG beads obtained from the MARTINI and SIRAH mappings (see SI, sec III). Our finding that the data-driven fit identifies an effective scale similar to that of these CG models thus supports the view that $\sim 4$ atoms per residue represents a natural balance point: it reduces complexity substantially while retaining the conformational ensemble characteristic of the atomistic system.
@@
The authors are indebted with M. Giulini for an insightful reading of the manuscript and useful comments. RP acknowledges support from ICSC - Centro Nazionale di Ricerca in HPC, Big Data and Quantum Computing, funded by the European Union under NextGenerationEU. Views and opinions expressed are however those of the author(s) only and do not necessarily reflect those of the European Union or The European Research Executive Agency. Neither the European Union nor the granting authority can be held responsible for them.
@@
Additional methodological details, full parameter lists, and supplementary analyses including relevance-resolution curve processing, PROPRE outputs, covariance and mapping entropy analyses, and comparison with coarse-grained models (PDF).